# GO4Align: Group Optimization for Multi-Task Alignment

### v1

@@ Abstract
This paper proposes GO4Align, a multi-task optimization approach that tackles task imbalance by explicitly aligning the optimization across tasks. To achieve this, we design an adaptive group risk minimization strategy, compromising two crucial techniques in implementation: (i) dynamical group assignment, which clusters similar tasks based on task interactions; (ii) risk-guided group indicators, which exploit consistent task correlations with risk information from previous iterations. Comprehensive experimental results on diverse typical benchmarks demonstrate our method’s performance superiority with even lower computational costs.   [^1] Code will be released in https://github.com/autumn9999/GO4Align.git.
@@ 1 Introduction
Figure 1: Performance and computational efficiency evaluation for MTO methods evaluated on NYUv2. Each method’s runtime is relative to a baseline method, which minimizes the sum of task-specific empirical risks. Left-down marks the comprehensive optimal results.
@@
Multi-task learning is a promising paradigm for handling several tasks simultaneously using a unified architecture. It can achieve data efficiency, improve generalization, and bring a reduction in computation cost compared with addressing each task individually Vandenhende et al. (2021). Due to these benefits, there is a growing surge of applications with multi-task learning in several domains, e.g., natural language processing Chen et al. (2021); Pilault et al. (2020); Liu et al. (2017), computer vision Vandenhende et al. (2021); Kendall et al. (2018); Liu et al. (2019) and reinforcement learning Espeholt et al. (2018); Sodhani et al. (2021). The crux of multi-task learning is to enable positive transfer among tasks while avoiding negative transfer, which usually exists among irrelevant tasks Xin et al. (2022); Liu et al. (2022); Standley et al. (2020).
@@
Existing Challenges: In avoiding the negative transfer, numerous multi-task optimization (MTO) methods Yu et al. (2020); Kendall et al. (2018); Navon et al. (2022); Xin et al. (2022); Liu et al. (2023) have emerged and attracted rising attention in recent years. A lasting concern in MTO is the task imbalance issue. It describes a phenomenon where some tasks are severely under-optimized Vandenhende et al. (2021), which can lead to worse overall performance with larger convergence differences across tasks.
@@
To deal with the task imbalance issue, scaling methods are proposed for MTO. According to differences in scaling manipulations, we roughly divide MTO methods into gradient-oriented Yu et al. (2020); Senushkin et al. (2023); Liu et al. (2021); Chen et al. (2020); Sener & Koltun (2018) and loss-oriented Kendall et al. (2018); Liu et al. (2019); Kurin et al. (2022); Lin et al. (2021b). The former tends to exhibit impressive results at the expense of higher computational or memory requirements during training time, due to the assessment of per-task gradients. In contrast, the latter preserves training-time efficiency but usually suffers from unsatisfactory overall performance. As shown in Fig. 1, most existing methods cannot simultaneously achieve superior performance and computational efficiency.
@@
Proposed Solution: To improve the overall performance and maintain computational and memory efficiency, we propose Group Optimization for multi-task Alignment (GO4Align), a novel and effective loss-oriented method in MTO. As shown in Fig. 2, this work identifies multi-task alignment as a crucial factor in solving task imbalance, which means learning progress across tasks should synchronously achieve superior performance over all tasks. GO4Align dynamically aligns learning progress across tasks by exploiting group-based task interactions for multi-task empirical risk minimization. The rationale behind this idea is that groupings can implicitly capture task correlations for more effective multi-task alignment and thus help multi-task optimizers benefit from positive interactions among relevant tasks. The primary contribution is two-fold:
@@
  • As a new member of the loss-oriented branch, GO4Align recasts the task imbalance issue to a bi-level optimization problem, yielding an adaptive group risk minimization principle for MTO. Such a principle allocates weights over task losses at a group level to achieve learning progress alignment among relevant tasks.
@@
  • We develop a heuristic optimization pipeline in GO4Align to tractably achieve the principle, involving dynamical group assignment and risk-guided group indicators. The pipeline incorporates beneficial task interactions into the group assignments and exploits task correlations for multi-task alignment, improving the overall multi-task performance.
@@
Experimental results show that our approach can outperform most existing state-of-the-art baselines in extensive benchmarks. Moreover, it does not sacrifice computational efficiency.
@@ 2 Preliminary
where ${\Omega}(\rm{\lambda})$ is a regularization term over task weights, designed to prevent the rapid collapse of these weights to zero, as discussed in Kendall et al. (2018). When all task weights are the same in scale, Scale-ERM will degenerate to the most simple strategy in MTO, where each task is treated equally during the joint training.
@@ 3 Methodology
In resolving the task imbalance issue effectively and efficiently, we develop GO4Align in this section. Such an approach relies on grouping-based task interactions to align learning progress across tasks, which are specific to the loss-oriented branch to hold the advantage of computational efficiency.
@@
Motivation of Multi-Task Alignment. Empirically, we can observe that the overall performance is worse when there is a larger convergence difference. The convergence difference is measured by the standard deviation of the task-specific epoch numbers to arrive at convergence. As shown in Figure 2, UW Kendall et al. (2018) underperforms FAMO Liu et al. (2023) in terms of a overall MTL metric $\Delta m\%$ (lower is better); while UW has a larger convergence difference than FAMO. Intuitively, a larger convergence difference means that per-task training dynamics are more asynchronous, usually leading to worse overall performance.
To perform MTO, we consider aligning tasks with group information in the multi-task risk minimization. We first present an adaptive group risk minimization principle for MTO, which targets the alignment of tasks’ learning progress from grouping-based task interactions. Then, in tractable problem-solving, we decompose the whole optimization process into two entangled phases: (i) dynamical group assignment and (ii) risk-guided group indicators. The pseudo-code of GO4Align is provided in Appendix B.
@@ 3.1 Adaptive Group Risk Minimization Principle
Recent advances Fifty et al. (2021); Kang et al. (2011) have explored incorporating multi-task grouping in feature sharing and revealed its benefit of aligning the learning progress through task interactions. Nevertheless, their grouping mechanisms ignore monitoring the learning progress, e.g., failure to capture variations in loss scales among tasks, weakening the effectiveness of multi-task alignment.
As a result, we propose taking task-specific learning dynamics, which directly impact the convergence behaviors in optimization, into grouping mechanism design. This induces the adaptive group risk minimization principle suitable for multi-task alignment. The hypothesis is that the dynamical grouping tends to implicitly exploit task correlations Fifty et al. (2021) and encourages beneficial task interactions from empirical risk information along the learning progress. Meanwhile, such a principle retains computational efficiency as it avoids the computations of per-task gradients.
Adaptive Group Risk Minimization (AGRM). We achieve beneficial task interactions by producing task weights with a grouping mechanism, which is adaptive to various loss scales and their learning dynamics over time. Then, we recast the task imbalance issue with the grouping mechanism into a bi-level optimization problem: (i) In the lower-level optimization, the model aims to cluster $M$ tasks of interest into $K$ groups. This implicitly exploits task correlations, where similar tasks should be clustered into one group, yielding more beneficial task interactions. With group assignments, group weights are designed to conduct learning progress alignment at the group level. (ii) In the upper-level optimization, the proposed principle updates model parameters from the grouped empirical risks, which implicitly relies on the lower-level optimized results. We illustrate such a bi-level optimization with grouping-based task interactions in Fig. 3.
@@
To be specific, we perform the bi-level optimization in Eq.(2) as the following steps. For the $t$-th iteration, we first compute the group information $\bm{\omega}_{t}$ and $\bm{\mathcal{G}}_{t}$ in the lower-level optimization; and then we update the model’s parameter in the upper-level optimization. As a result, we obtain the updated parameter $\bm{\theta}_{t+1}$, which is used to compute task-specific risk information at the next iteration. In the proposed principle, intra-group tasks share the same scaling weight $\omega^{k}\geq 0$ to prevent similar tasks from inconsistent learning progress, improving knowledge sharing among similar tasks.
Important in AGRM is to accommodate the grouping during training dynamically. In practice, the proposed principle is compatible with any gradient-based optimizer, such as SGD and Adam Kingma & Ba (2014), yielding dynamical training for each task. As a new member of the loss-oriented branch, the grouping assignment matrix and group weights in GO4Align can sufficiently utilize the loss information over time to adaptively assign tasks and weight groups.
Unlike prior works on multi-task grouping Fifty et al. (2021); Kang et al. (2011), which require group-specific architectures, our proposed principle focuses on group-specific scaling and adaptively executing grouping operations. These insights are: (i) the criteria for grouping tasks should take both learning dynamics and loss scales into consideration so that similar tasks can benefit from each other’s intermediate feature information and boost performance; (ii) adaptive grouping aligns learning progress across tasks and provides a more effective way for across-task information transfer.
@@ 3.2 Dynamical Group Assignment
Intuitively, we conduct the group assignment as a clustering process based on these group indicators $\bm{\gamma}_{t}(\bm{\theta}_{t})$. In this case, each cluster represents a group, and the cluster center corresponds to the corresponding group weight. There exist numerous clustering algorithms available to achieve this, and this work takes the K-means clustering algorithm Krishna & Murty (1999); Kodinariya et al. (2013); Ahmed et al. (2020) as a practical clustering implementation. Next, we can specify the optimization objective of the dynamical group assignment as:
@@
where $\bm{\omega}_{t}^{\top}={\bm{\gamma}_{t}^{\top}}(\bm{\theta}_{t})\bm{\mathcal{G}}_{t}^{-1}$ indicates that the cluster center closely relates to the group assignment matrix. The designed dynamical group assignment plays an important role in the lower-level optimization of the proposed AGRM, and it tends to cluster similar tasks into the same group while scattering dissimilar ones in clusters.
By integrating the dynamical group assignment in Eq. (3) and the group indicators into Eq. (2), we can provide an instantiation for the AGRM’s optimization objective:
@@
Moreover, the dynamic group assignment heuristically clusters tasks from the group indicators, which avoids the exhausted search of appropriate task combinations for performance gains like previous works  Fifty et al. (2021).
@@ 3.3 Risk-guided Group Indicators
The misalignment of learning across tasks can usually be attributed to various risk scales and asynchronous learning dynamics among tasks over time. To address this, we take two operations, scale-balance and smooth-alignment, into consideration and then obtain the risk-guided group indicators by combining them.
@@
where $\rm{diag}(\cdot)$ constructs a diagonal matrix with the elements of the vector placed on the diagonal. $\bar{\hat{\bm{L}}}(\bm{\theta}_{t})$ is a scalar to represent the average risk, and $[\cdot]_{M}$ represents the construction of an $M$-dimentional vector whose elements are all equal to the average risk. To avoid the upper-level optimization degenerating into a fixed scalar, the gradients of empirical risks in the lower-level optimization are not being computed through them. However, in practice, the learning dynamics over time tend to make the scale vector inconsistent over iterations Liu et al. (2021); Yu et al. (2020), which is not conducive to aligning learning progress. We therefore introduce smooth-alignment to update the scale vector with historical information from previous iterations.
Smooth-alignment. To avoid sudden fluctuations of scale vectors over iterations, we introduce the smoothness vector $\mathcal{Q}_{t}(\bm{\theta}_{1:t})=\left[q_{t}^{1}(\bm{\theta}_{1:t}),q_{t}^{2}(\bm{\theta}_{1:t}),\dots,q_{t}^{M}(\bm{\theta}_{1:t})\right]^{\top}\in\mathbb{R}^{M}$, which smooths the updating of the scale vector with previous risk information. Thus, the smoothness vector can deal with asynchronous learning dynamics over time, which helps the model to reduce the imbalance of training across tasks. To be specific, we compute the smoothness vector by a normalized exponential moving average as follows:
$\mathcal{Q}_{t}(\bm{\theta}_{1:t})$ $=\sigma\Big{[}\bm{\mathcal{Q}}_{t-1}(\bm{\theta}_{1:t-1})\odot\exp(-\beta\hat{\bm{L}}(\bm{\theta}_{t}))\Big{]},$ (6)
where $\odot$ denotes the element-wise multiplication and $\sigma[\cdot]$ normalizes the sum of all smoothness elements to be $1$. $\beta$ is a temperature hyperparameter to control the influence of previous grouping information. Note that when $\beta$ is close to zero, each element in the smoothness vector will degrade to a fixed value $\frac{1}{M}$, which does not capture any historical information to group indicators.
Risk-guided Group indicators. By integrating Eq. (5) and (6), we obtain the group indicators by:
@@
Then, based on the group indicators $\bm{\gamma}_{t}(\bm{\theta}_{t})$ with scale-balance and smooth-alignment from risk information, we optimize the dynamical group assignment as Eq.(3) to assign tasks into groups in the lower-level optimization.

### v2

@@ Abstract
This paper proposes GO4Align, a multi-task optimization approach that tackles task imbalance by explicitly aligning the optimization across tasks. To achieve this, we design an adaptive group risk minimization strategy, comprising two techniques in implementation: (i) dynamical group assignment, which clusters similar tasks based on task interactions; (ii) risk-guided group indicators, which exploit consistent task correlations with risk information from previous iterations. Comprehensive experimental results on diverse benchmarks demonstrate our method’s performance superiority with even lower computational costs.
@@ 1 Introduction
Figure 1: Performance and computational efficiency evaluation for MTO methods evaluated on NYUv2. Each method’s training time is relative to a baseline method, which minimizes the sum of task-specific empirical risks. Left-bottom marks comprehensive optimal results.
@@
Multi-task learning is a promising paradigm for handling several tasks simultaneously using a unified architecture. It can achieve data efficiency, improve generalization, and reduce computation costs compared with addressing each task individually [1]. Due to these benefits, there is a growing surge of applications with multi-task learning in several domains, e.g., natural language processing [2, 3, 4], computer vision [1, 5, 6] and reinforcement learning [7, 8]. The crux of multi-task learning is to enable positive transfer among tasks while avoiding negative transfer, which usually exists among irrelevant tasks [9, 10, 11].
@@
Existing Challenges: In avoiding the negative transfer, numerous multi-task optimization (MTO) methods [12, 5, 13, 9, 14] have emerged and attracted rising attention in recent years. A lasting concern in MTO is the task imbalance issue. It describes a phenomenon where some tasks are severely under-optimized [1], which can lead to worse overall performance with larger convergence differences across tasks.
@@
To deal with the task imbalance issue, scaling methods are proposed for MTO. According to differences in scaling manipulations, we roughly divide MTO methods into gradient-oriented [12, 15, 16, 17, 18] and loss-oriented [5, 6, 19, 20]. The former tends to exhibit impressive results at the expense of higher computational or memory requirements during training time due to the assessment of per-task gradients [19]. [^1] In MTO, computational efficiency refers to training time efficiency. In contrast, the latter preserves training-time efficiency but usually suffers from unsatisfactory overall performance. As shown in Fig. 1, most existing methods cannot simultaneously achieve superior performance and computational efficiency.
@@
Proposed Solution: To improve the overall performance and maintain computational and memory efficiency, we propose Group Optimization for multi-task Alignment (GO4Align), a novel and effective loss-oriented method in MTO. As shown in Fig. 2, this work identifies multi-task alignment as a crucial factor in solving task imbalance, which means learning progress across tasks should synchronously achieve superior performance over all tasks. The proposed model dynamically aligns learning progress across tasks by exploiting group-based task interactions for multi-task empirical risk minimization. The rationale behind this is that groupings can implicitly capture task correlations for more effective multi-task alignment and thus help multi-task optimizers benefit from positive interactions among relevant tasks. The primary contribution is two-fold:
@@
  • As a new member of the loss-oriented MTO branch, GO4Align recasts the task imbalance issue to a bi-level optimization problem, yielding an adaptive group risk minimization principle for MTO. Such a principle allocates weights over task losses at a group level to achieve learning progress alignment among relevant tasks.
@@
  • We develop a heuristic optimization pipeline in GO4Align to tractably achieve the principle, involving dynamical group assignment and risk-guided group indicators. The pipeline incorporates beneficial task interactions into the group assignments and exploits task correlations for multi-task alignment, improving overall multi-task performance.
@@
Experimental results show that our approach can outperform existing state-of-the-art baselines in extensive benchmarks. Moreover, it does not sacrifice computational efficiency.
@@ 2 Preliminary
where ${\Omega}(\bm{\lambda})$ is a regularization term over task weights, designed to prevent the rapid collapse of these weights to zero, as discussed in Kendall et al. [5]. When all task weights are the same in scale, Scale-ERM will degenerate to the most simple strategy in MTO, where each task is treated equally during the joint training.
@@ 3 Methodology
In resolving the task imbalance issue effectively and efficiently, we develop GO4Align in this section. Our approach relies on grouping-based task interactions to align learning progress across tasks. As a new member of the loss-oriented branch, GO4Align holds the advantage of computational efficiency without the requirements of per-task gradients.
@@
Motivation of Multi-Task Alignment. Empirically, we can observe that the overall performance is worse when there is a larger convergence difference between tasks. The convergence difference is measured by the standard deviation of the task-specific epoch numbers to reach convergence. As shown in Fig. 2, UW [5] underperforms FAMO [14] in terms of a overall MTL metric $\Delta m\%$ (lower is better); while UW has a larger convergence difference than FAMO. Intuitively, a larger convergence difference means that per-task training dynamics are more asynchronous, usually leading to worse overall performance.
To perform MTO, we consider aligning tasks with group information in the multi-task risk minimization. We first present an adaptive group risk minimization principle for MTO, which targets the alignment of tasks’ learning progress from grouping-based task interactions. Then, in tractable problem-solving, we decompose the whole optimization process into two entangled phases: (i) dynamical group assignment and (ii) risk-guided group indicators. The pseudo-code of GO4Align is provided in Appendix A.
@@ 3.1 Adaptive Group Risk Minimization Principle
Recent advances [21, 22] have explored incorporating multi-task grouping in feature sharing and revealed its benefit of aligning the learning progress through task interactions. Nevertheless, their grouping mechanisms ignore monitoring the learning progress, e.g., failure to capture variations in loss scales among tasks, weakening the effectiveness of multi-task alignment.
As a result, we design a new grouping mechanism for MTO with task-specific learning dynamics, which directly impacts the convergence behaviors in optimization. This induces the adaptive group risk minimization principle suitable for multi-task alignment. The hypothesis is that the dynamical grouping tends to implicitly exploit task correlations [21] and encourages beneficial task interactions from empirical risk information along the learning progress. Meanwhile, such a principle retains computational efficiency as it avoids the computations of per-task gradients.
Adaptive Group Risk Minimization (AGRM). We first achieve beneficial task interactions by producing task weights with a grouping mechanism, which is adaptive to various loss scales and their learning dynamics over time. Then, we recast the task imbalance issue with the grouping mechanism into a bi-level optimization problem: (i) In the lower-level optimization, the model aims to cluster $M$ tasks of interest into $K$ groups. This implicitly exploits task correlations, where similar tasks should be clustered into one group, yielding more beneficial task interactions. With group assignments, group weights are designed to conduct learning progress alignment at the group level. The group weights in the multi-task objective are motivated as an extension of the observation that similar tasks benefit greatly from training together through parameter sharing [21]. (ii) In the upper-level optimization, the proposed principle updates model parameters from the grouped empirical risks, which implicitly relies on the lower-level optimized results. We illustrate GO4Align’s bi-level optimization with grouping-based task interactions in Fig. 3.
@@
To be specific, we perform the bi-level optimization in Eq. (2) according to the following steps. For the $t$-th iteration, we first compute the group information $\bm{\omega}_{t}$ and $\bm{\mathcal{G}}_{t}$ in the lower-level optimization; and then we update the model’s parameter in the upper-level optimization. As a result, we obtain the updated parameter $\bm{\theta}_{t+1}$, which is used to compute task-specific risk information at the next iteration. In the proposed principle, intra-group tasks share the same scaling weight $\omega^{k}\geq 0$ to prevent similar tasks from inconsistent learning progress, improving knowledge sharing among similar tasks.
Important in AGRM is to accommodate the grouping during training dynamically. In practice, the proposed principle is compatible with any gradient-based optimizer, such as SGD and Adam [23], yielding dynamical training for each task. As a new member of the loss-oriented branch, the grouping assignment matrix and group weights in GO4Align can sufficiently utilize the loss information over time to adaptively assign tasks and weight groups.
Unlike prior works on multi-task grouping [21, 22], which require group-specific architectures, our proposed principle focuses on group-specific scaling and adaptively executing grouping operations. Considering the differences in architecture and optimization within grouping mechanisms, we provide two insights: (i) the criteria for grouping tasks should take both learning dynamics and loss scales into consideration so that similar tasks can benefit from each other’s intermediate feature information and boost performance; (ii) adaptive grouping aligns learning progress across tasks and provides a more effective way for across-task information transfer.
@@ 3.2 Dynamical Group Assignment
Intuitively, we conduct the group assignment as a clustering process based on these group indicators $\bm{\gamma}_{t}(\bm{\theta}_{t})$. In this case, each cluster represents a group, and the cluster center is set to the group weight. Many clustering algorithms are available to achieve this. In this work, we take the K-means clustering algorithm [24, 25, 26] as a practical clustering implementation. Thus, we specify the optimization objective of the dynamical group assignment as:
@@
where $\bm{\omega}_{t}^{\top}={\bm{\gamma}_{t}^{\top}}(\bm{\theta}_{t})\bm{\mathcal{G}}_{t}^{-1}$ indicates that the cluster center closely relates to the group assignment matrix. It is worth noting that $\bm{\mathcal{G}}_{t}^{-1}$ is a generalized inverse, especially one-sided right inverse, $\bm{\mathcal{G}}^{-1}=\bm{\mathcal{G}}^{\top}(\bm{\mathcal{G}}\bm{\mathcal{G}}^{\top})^{-1}.$ The designed dynamical group assignment plays an important role in the lower-level optimization of the proposed AGRM, and it tends to cluster similar tasks into the same group while scattering dissimilar ones in clusters.
By integrating the dynamical group assignment in Eq. (3) and the group indicators into Eq. (2), we provide an instantiation for the AGRM’s optimization objective:
@@
Moreover, the dynamic group assignment heuristically clusters tasks from the group indicators, which avoids the exhausted search of appropriate task combinations for performance gains like previous works  [21].
@@ 3.3 Risk-guided Group Indicators
The misalignment of learning across tasks can usually be attributed to various risk scales and asynchronous learning dynamics among tasks over time. To address this, we take two operations with risk information, scale-balance and smooth-alignment, into consideration and then obtain the risk-guided group indicators by combining them. The role of the group indicators is to use the risk information to explore the relationships among tasks without incurring the expensive computational cost associated with gradients. Compared with other loss-oriented methods, our group indicator can capture the differences in the per-task risk scale and fully utilize the learning dynamics over time, yielding better representations of risk information.
@@
where $\rm{diag}(\cdot)$ constructs a diagonal matrix with the elements of the vector placed on the diagonal. $\bar{\hat{\bm{L}}}(\bm{\theta}_{t})$ is a scalar to represent the average risk, and $[\cdot]_{M}$ represents the construction of an $M$-dimentional vector whose elements are all equal to the average risk. To avoid the upper-level optimization degenerating into a fixed scalar, the gradients of empirical risks in the lower-level optimization are not being computed through them. However, in practice, the learning dynamics over time tend to make the scale vector inconsistent over iterations [16, 12], which is not conducive to aligning learning progress. We therefore introduce smooth-alignment to update the scale vector with historical information from previous iterations.
Smooth-alignment. To avoid sudden fluctuations of scale vectors over iterations, we introduce the smoothness vector $\mathcal{Q}_{t}(\bm{\theta}_{1:t})=\left[q_{t}^{1}(\bm{\theta}_{1:t}),q_{t}^{2}(\bm{\theta}_{1:t}),\dots,q_{t}^{M}(\bm{\theta}_{1:t})\right]^{\top}\in\mathbb{R}^{M}$, which smooths the updating of the scale vector with previous risk information. Thus, the smoothness vector can deal with asynchronous learning dynamics over time, which helps the model reduce the imbalance of training across tasks. To be specific, we compute the smoothness vector by a normalized exponential moving average as follows:
$\mathcal{Q}_{t}(\bm{\theta}_{1:t})$ $=\sigma\Big{[}{\mathcal{Q}}_{t-1}(\bm{\theta}_{1:t-1})\odot\exp(-\beta\hat{\bm{L}}(\bm{\theta}_{t}))\Big{]},$ (6)
where $\odot$ denotes the element-wise multiplication and $\sigma[\cdot]$ normalizes the sum of all smoothness elements to be $1$. $\beta$ is a temperature hyperparameter to control the influence of current risk information. Note that when $\beta$ is close to zero, each element in the smoothness vector will degrade to a fixed value $\frac{1}{M}$, which does not capture any historical information to group indicators.
Risk-guided Group Indicators. By element-wise multiplying the scale vector in Eq. (5) and the smoothness vector in Eq. (6), we obtain the group indicators with sufficient risk information as:
@@
Then, with the group indicators $\bm{\gamma}_{t}(\bm{\theta}_{t})$, we optimize the dynamical group assignment in Eq. (3) to assign tasks into groups in the lower-level optimization.